It took 3 hours to row a boat 18 km against the current. The return trip with the current took 1.5 hours. Find the speed of the rowboat in still water.
Answer by Paul(988)

3(t-18)=1.5(t+18)
3t-54=1.5t+27
3t=1.5t+81
1.5t=81
t=54
Therefore, the speed of the boat in still water is 54km